Find three consecutive odd integers such that the sum of twice the first, twice the second, and three times the third is 9. An explanation would be great too!
Found 2 solutions by ewatrrr, stanbon:
Answer by ewatrrr(24785)


Hi
consecutive odd integers differ by two
Let x, (x+2),(x+4) represent the three consecutive odd integers
Question states***Write as we Read
2x + 2(x+2) + 3(x+4) = 9
solving for x
2x + 2x+4 + 3x+12 = 9
7x + 16 = 9
7x = -7
x = -1 the three consecutive odd integers are: -1, 1, 3
CHECKING our Answer***
-2 + 2 + 9 = 9

Answer by stanbon(75887)

Find three consecutive odd integers such that the sum of twice the first, twice the second, and three times the third is 9.
---
Odd numbers must be represented as an odd # more than or less than
an even number.
-----
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
---------------
Equation:
2(2x-1)+2(2x+1)+3(2x+3) = 9
4x-2+4x+2+6x+9 = 9
14x + 9 = 9
14x = 0
x = 0
---
1st: -1
2nd: +1
3rd: 3
---
Check:
2(-1)+2(1)+3(3) = 9
checks
==============
Cheers,
Stan H.